OpenStudy (anonymous):

solve: (w+1)^(3/2)=64

OpenStudy (anonymous):

Or you could just post another question. Haha. w+1=64^(2/3) w+1=16 w=15. I'd appreciate it if you fanned me.

OpenStudy (anonymous):

ye i would but i have no idea how to

OpenStudy (anonymous):

Click "Become a fan" next to my name if you haven't already.

OpenStudy (anonymous):

i dont c were it says that it only say superhero(39fans)

OpenStudy (anonymous):

Huh. Oh well. Haha.

OpenStudy (anonymous):

omg 3 questions lefrt!@@@@

OpenStudy (anonymous):

For what? lol